An engineer designed a valve that will regulate water pressure on an automobile engine. The engineer designed the valve such that it would produce a mean pressure of 4.5 pounds/square inch. It is believed that the valve performs above the specifications. The valve was tested on 200 engines and the mean pressure was 4.6 pounds/square inch. Assume the standard deviation is known to be 0.6. A level of significance of 0.02 will be used. Find the value of the test statistic. Round your answer to 2 decimal places.
Answer #1

This is the two tailed test .

The null and alternative hypothesis is

H0 :  \mu =

Ha : \mu \neq

Test statistic = z

= (\bar x - \mu ) / \sigma / \sqrt n

= (4.6 - 4.5) / 0.6 / \sqrt 200

Test statistic = 2.36

P(z > 2.36) = 1 - P(z < 2.36) = 0.0091

P-value = 0.0182